OpenStudy (alesia):

can you help me find the first derivative of y=square root of cos(X) + cox (square root of x) ?

OpenStudy (anonymous):

\(\sqrt{cos{x}}+cos{\sqrt{x}}\) ?

OpenStudy (anonymous):

I never remember the rule for deriving square roots, so instead I tend to rewrite roots as exponent and then just use exponent rule. \((cos{x})^{\frac{1}{2}} + cos{(x^{\frac{1}{2}})}\) Use chain rule, exponent rule, and the rule for derivative of cosine.